# references.bib

### v4

@@ 1.4 Heisenberg calculus for differential operators and H-elliptic operators
$P$ is of H-order $\leq k$ if $P$ can be locally expressed as a sum of monomials of the form
@@
$\sigma_{H}^{k}(P,x,\pi)=\sum_{\text{order}=k}d\pi([X_{1}]_{x})\cdots d\pi([X_{s}]_{x})\otimes\Phi_{x}:\mathcal{H}^{\infty}_{\pi}\otimes E_{x}\to\mathcal{H}^{\infty}_{\pi}\otimes F_{x},$
@@
  • The $KK$-classes we will define later and the resulting map $\sigma_{H}^{k}(P,x,\pi)$ depend only on the principal component of $P$, so we can select the bundle map $TM\to gr(H)$ to be the simplest one which maps $v\in H^{i}_{x}$ to $[v]\in H^{i}_{x}/H^{i-1}_{x}\subset gr(H)_{x}$.
@@
Definition 1.4.5.
Let $P=\nabla_{X_{1}}\cdots\nabla_{X_{s}}\Phi$, then
  • At each point $x\in M$, we define
    as the freezing coefficient of $P$ at $x$,
  • The H-cosymbol of $P$ is defined to be the family of operators
    $\sigma_{H}(P):=\{\sigma_{H}(P,x)\}_{x\in M}=\{P_{x}\}_{x\in M}.$
    on $\{Gr(H)_{x}\}$ parameterized by $x\in M$.
The Heisenberg principal cosymbol of $P$ at $\pi$ $\sigma_{H}^{k}(P,x,\pi)$ can be calculated by
Remark 1.4.6.
$\sigma_{H}(P)$ defined above is different from the symbol in the ordinary sense. For an ordinary differential operator of order $m$, $P=\sum_{|\alpha|\leq m}a_{\alpha}\partial_{x}^{\alpha}$, freezing the coefficient at a point $x\in M$ induces the constant coefficient operator $P_{x}=\sum_{|\alpha|=m}a_{\alpha}(x)\partial_{\xi}^{\alpha}$ on $T_{x}M$. This family $\{P_{x}\}_{x\in M}$ is called the ordinary principal “cosymbol” of $P$, with its Fourier transform being the principal symbol of $P$ in the ordinary sense:
Let $N=\dim Gr(H)$. Let $\{X_{j}\}^{N}_{j=1}$ represent a local basis of the vector bundle $gr(H)$ near $x\in M$. then, $\{Y_{j}=X_{j}(x)\}^{N}_{j=1}$ will be a basis of $gr(H)_{x}$ at each point near $x$. Each $Y_{j}\in gr(H)_{x}$ can be regarded as a first-order left-invariant differential operator on $Gr(H)_{x}$, the formal sum
@@
For each $x\in M$, since $\sigma_{H}(P,x)$ corresponds to a left-invariant differential operator on $Gr(H)_{x}$, it follows that for $f,g\in C_{c}^{\infty}(Gr(H)_{x})$, $P(f\ast g)=Pf\ast g$. This principle allows us to extend the operator to the group $C^{\ast}$-algebra $C^{\ast}(Gr(H)_{x})$ and yields
@@
This can be considered to be a morphism between $C^{\ast}(Gr(H)_{x})$-Hilbert modules. Thus, the cosymbol $\sigma_{H}(P)=\{\sigma_{H}(P,x)\}$ parametrized by $x\in M$ can be extended to an unbounded multiplier of $C^{\ast}({Gr(H)})$
$\sigma_{H}(P):C(M;E)\otimes_{C(M)}C^{\ast}(Gr(H))\to C(M;F)\otimes_{C(M)}C^{\ast}(Gr(H)).$
The representation $\pi$ can be extended to a representation of $C^{\ast}(Gr(H)_{x})$ on $\mathcal{H_{\pi}}$, thus $\sigma_{H}(P)\otimes_{\pi}1$ on $(E_{x}\otimes C^{\ast}(Gr(H)_{x}))\otimes_{\pi}\mathcal{H_{\pi}}=E_{x}\otimes\mathcal{H_{\pi}}$ yields an operator
@@
    The range and target map on the groupoid $Gr(H)$ denote the same map. Therefore, the groupoid $C^{\ast}$-algebra structure of $C^{\ast}(Gr(H))$ is a smooth family of group $C^{\ast}$-algebras $\{C^{\ast}{(Gr(H)_{x})}\}_{x\in M}$.
For left-invariant and homogeneous differential operators on a homogeneous Lie group, hypoellipticity is equivalent to the Rockland condition that we will define in the following theorem. The Rockland condition also aids in comprehending the definition of H-ellipticity.
Theorem 1.4.9.
[Rockland theorem, [melin1983parametrix]] The principal cosymbol $\sigma_{H}(P,x)$ of a differential operator $P$ at $x\in M$ is left invertible on $L^{2}(M)$ if and only if for every unitary non-trivial representation $\pi$ of the group $Gr(H)_{x}$, the (unbounded) operator $\pi(\sigma_{H}(P,x))$ is injective.
@@ 1.5 Hilbert modules and $KK$-theory
In this section, we briefly introduce the Hilbert module theory and Kasparov’s $KK$-theory. See [blackadar1998k] for details.
@@
Let $A$, $B$ be separable and $D$ be $\sigma$-unital graded $C^{\ast}$-algebras, let $(E_{1},\phi_{1},F_{1})$ be a Kasparov $(A,D)$-module and $(E_{2},\phi_{2},F_{2})$ be a Kasparov $(D,B)$-module, there is a Kasparov product for $(E_{1},\phi_{1},F_{1})$ and $(E_{2},\phi_{2},F_{2})$, which is a Kasparov $(A,B)$-module, unique up to homotopy $\sim_{h}$. The Kasparov product defines a bilinear function ${\otimes}_{D}:KK(A,D)\times KK(D,B)\rightarrow KK(A,B)$.
@@ 1.6 Heisenberg calculus for pseudo-differential operators
The normalization of the operators and inversion of cosymbol classes modulo smoothing terms are essential in our theory, but they are outside of the class of differential operators. In this section, we consider pseudo-differential operators.
@@
A subset $X$ of a groupoid $\mathcal{G}$ is proper if both $r:G\rightarrow M$ and $s:X\rightarrow M$ are proper maps. Denote by $\Omega_{r}$ and $\Omega_{s}$ the bundle of smooth densities tangent to the range and source fibers of $\mathcal{G}$ respectively. The space of smooth sections of $\Omega_{r}$ whose support is a proper subset of $\mathcal{G}$ is denoted $C_{p}^{\infty}(\mathcal{G};\Omega_{r})$.
@@
$\mathbb{G}r(H)$ is defined to be the disjoint union of the osculating groupoid $Gr(H)$ with a family of pair groupoids indexed by ×=$\mathbb{R}^{\times}=\mathbb{R}\setminus\{0\}$,
@@
it is called the H-tangent groupoid of the osculating groupoid $Gr(H)$. The main result of [van2017tangent] is that there is a compatible global smooth structure on $\mathbb{G}r(H)$ that makes $\mathbb{G}r(H)$ into a Lie groupoid.
@@
The cosymbol of a pseudo-differential operator can be viewed as a family of operators on the manifolds $Gr(H)_{x}$. So specifically, we need to analyze the case special case $\mathcal{G}=Gr(H)$.
@@
Let $S^{m}(\mathcal{G})$ be the space of distributions $u\in\mathcal{E}^{\prime}(\mathcal{G})$ such that $u$ is transversal to the map $\pi:\mathcal{G}\to M$ and for each $\lambda\in\mathbb{R}^{+}$,
${\alpha_{\lambda}}_{\ast}u-\lambda^{m}u\in C_{c}^{\infty}(\mathcal{G}).$
Let $u\in S^{m}(\mathcal{G})$. For each $x\in M$, the operator can be defined as
@@
Let $u\in S^{m}(\mathcal{G})$, then
  • If $m\leq 0$, then $\overline{Op(u)}\in\mathcal{L}(C^{\ast}(\mathcal{G}))$;
  • If $m<0$, then $\overline{Op(u)}\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$;
  • If $m>0$, and $u$ satisfy the Rockland condition, then $\overline{Op(u)}$ is a regular operator.
For left-invariant differential operators on $\mathcal{G}$, the Rockland condition facilitates the consideration of the $KK$-class:
@@
$D=\overline{P}/(1+\overline{P}^{\ast}\overline{P})^{1/2}\in\mathcal{L}(C(M;E)\otimes_{C(M)}C^{\ast}(\mathcal{G}),C(M;F)\otimes_{C(M)}C^{\ast}(\mathcal{G}))$
@@
For any arbitrary H-elliptic operator $P$, the principal cosymbol $\sigma_{H}(P)$ represents a $KK$-element $[\sigma_{H}(P)]\in KK(\mathbb{C},C^{\ast}(Gr(H))\cong K_{0}(C^{\ast}(Gr(H)))$.
@@
$\Big{[}C(M;E_{0}\oplus E_{1})\otimes_{C(M)}C^{\ast}(\mathcal{G}),\begin{pmatrix}0&\overline{P}^{\ast}/(1+\overline{P}\overline{P}^{\ast})^{1/2}\\ \overline{P}/(1+\overline{P}^{\ast}\overline{P})^{1/2}&0\end{pmatrix}\Big{]}$
represents an element in $KK(\mathbb{C},C^{\ast}(\mathcal{G}))\cong K_{0}(C^{\ast}(\mathcal{G}))$. Apply this result to $\sigma_{H}(P)$, the family of left invariant differential operators on $Gr(H)$, it represents an element in $KK(\mathbb{C},C^{\ast}(Gr(H))\cong K_{0}(C^{\ast}(Gr(H)))$. ∎
Later on we identify $KK(\mathbb{C},A)$ with $K_{0}(A)$ and $KK(A,\mathbb{C})$ with $K^{0}(A)$.
This framework can be extended to pseudo-differential operators [androulidakis2022pseudodifferential]. For a pseudo-differential operator $P$ on $M$, the principal cosymbol $\sigma_{H}(P)$ is an unbounded multiplier of the $C^{\ast}$-algebra $C^{\ast}(Gr(H))$, and the H-order of $P$ less or equal to 00 decide the boundedness or compactness of $P$ as follows:
@@
Let $P$ be a pseudo-differential operators of H-order $k$ on $M$, then
@@
On a manifold $M$ with the filtration structure $H^{\bullet}$, a sequence of Sobolev spaces $\tilde{H}^{s}(M,E,F)$ are defined in [androulidakis2022pseudodifferential]. For a non-negative integer $s$, it encompasses the set of vector bundle $E$-valued distributions $u$ such that $Du\in L^{2}(F)$ for any differential operators $D\in\text{Diff}^{s}_{{H}}(M,E,F)$. It can be naturally extended to real numbers $s\geq 0$, and for $s<0$, $\tilde{H}^{s}(M)$ is defined to be the dual of $\tilde{H}^{-s}(M)$.

### v5

@@ 1.4 Heisenberg calculus for differential operators and H-elliptic operators
In this section, we will see how to use Kirillov’s orbit method in Section 1.2 to analyze the principal cosymbol of a differential operator on a filtered manifold $M$.
@@
We say $P$ is of H-order $\leq k$ if $P$ can be locally expressed as a sum of monomials of the form
@@
$\sigma_{H}^{k}(P,x,\pi)=\sum_{\text{H-order}=k}d\pi([X_{1}]_{x})\cdots d\pi([X_{s}]_{x})\otimes\Phi_{x}:\mathcal{H}^{\infty}_{\pi}\otimes E_{x}\to\mathcal{H}^{\infty}_{\pi}\otimes F_{x},$
@@
  • The $KK$-classes we will define later and the resulting map $\sigma_{H}^{k}(P,x,\pi)$ depend only on the H-principal component of $P$, so we can select the bundle map $TM\to gr(H)$ to be the simplest one which maps $v\in H^{i}_{x}$ to $[v]\in H^{i}_{x}/H^{i-1}_{x}\subset gr(H)_{x}$.
@@
Let $N=\dim Gr(H)$. Let $\{X_{j}\}^{N}_{j=1}$ represents a local basis of the vector bundle $gr(H)$ near $x\in M$. then, $\{Y_{j}=X_{j}(x)\}^{N}_{j=1}$ will be a basis of $gr(H)_{x}$ at each point near $x$. Each $Y_{j}\in gr(H)_{x}$ can be regarded as a first-order left-invariant differential operator on $Gr(H)_{x}$, the formal sum
@@
Definition 1.4.6.
Let $P=\nabla_{X_{1}}\cdots\nabla_{X_{s}}\Phi$, then
  • At each point $x\in M$, we define
    as the freezing coefficient of $P$ at $x$.
  • The H-cosymbol of $P$ is defined to be the family of operators
    $\sigma_{H}(P):=\{\sigma_{H}(P,x)\}_{x\in M}=\{P_{x}\}_{x\in M}.$
    on $\{Gr(H)_{x}\}$ parameterized by $x\in M$.
By Lemma 1.4.4 and the formula of $\sigma_{H}^{k}(P,x,\xi)$ in Definition 1.4.2, the Heisenberg principal cosymbol of $P$ at $\pi$ can be calculated as
Remark 1.4.7.
The Heisenberg principal cosymbol $\sigma_{H}(P)$ defined above is different from the symbol in the ordinary sense. For an ordinary differential operator of order $m$, $P=\sum_{|\alpha|\leq m}a_{\alpha}\partial_{x}^{\alpha}$, freezing the coefficient at a point $x\in M$ induces the constant coefficient operator $P_{x}=\sum_{|\alpha|=m}a_{\alpha}(x)\partial_{\xi}^{\alpha}$ on $T_{x}M$. This family $\{P_{x}\}_{x\in M}$ is called the ordinary principal “cosymbol” of $P$, with its Fourier transform being the principal symbol of $P$ in the ordinary sense:
@@
For each $x\in M$, since $\sigma_{H}(P,x)$ corresponds to a left-invariant differential operator on $Gr(H)_{x}$, it follows that for $f,g\in C_{c}^{\infty}(Gr(H)_{x})$, $\sigma_{H}(P,x)(f\ast g)=f\ast\sigma_{H}(P,x)(g)$. We can extend the operator to the group $C^{\ast}$-algebra $C^{\ast}(Gr(H)_{x})$ and yields
@@
This can be considered to be a morphism between $C^{\ast}(Gr(H)_{x})$-Hilbert modules. Thus, the cosymbol $\sigma_{H}(P)=\{\sigma_{H}(P,x)\}$ parametrized by $x\in M$ can be extended to an unbounded multiplier of $C^{\ast}({Gr(H)})$:
$\overline{\sigma_{H}(P)}:C_{0}(M;E)\otimes_{C_{0}(M)}C^{\ast}(Gr(H))\to C_{0}(M;F)\otimes_{C_{0}(M)}C^{\ast}(Gr(H)).$
The representation $\pi$ can be extended to a representation of $C^{\ast}(Gr(H)_{x})$ on $\mathcal{H_{\pi}}$, thus $\overline{\sigma_{H}(P)}\otimes_{\pi}1$ on $(E_{x}\otimes C^{\ast}(Gr(H)_{x}))\otimes_{\pi}\mathcal{H_{\pi}}=E_{x}\otimes\mathcal{H_{\pi}}$ yields an operator
@@
    The range and sourse map on the groupoid $Gr(H)$ are the same map. Therefore, the groupoid $C^{\ast}$-algebra structure of $C^{\ast}(Gr(H))$ is a smooth family of group $C^{\ast}$-algebras $\{C^{\ast}{(Gr(H)_{x})}\}_{x\in M}$.
For left-invariant and homogeneous differential operators on a homogeneous Lie group, hypoellipticity is equivalent to the Rockland condition that we will define in the following theorem. The Rockland condition also helps to understand the definition of H-ellipticity.
Theorem 1.4.9 (Rockland theorem, [melin1983parametrix]).
The principal cosymbol $\sigma_{H}(P,x)$ of a differential operator $P$ at $x\in M$ is left invertible on $L^{2}(M)$ if and only if for every unitary non-trivial representation $\pi$ of the group $Gr(H)_{x}$, the (unbounded) operator $\pi(\sigma_{H}(P,x))$ is injective.
@@ 1.5 Hilbert modules and $KK$-theory
In order to analyze the symbol class and index class of H-elliptic operators, we need to understand the $KK$-theory. In this section, we briefly introduce the Hilbert module theory and Kasparov’s $KK$-theory. See [blackadar1998k] for details.
@@
When $A$ and $B$ are graded $G$-algebras, the equivariant $KK$-group $KK^{G}(A,B)$ is defined in a similar manner with the additional condition that $F$ is $G$-continuous and $(g\cdot F-F)\phi(a)\in\mathcal{K}(E)$ for all $a\in A$ and $g\in G$.
Example 1.5.9.
In the Kasparov $K$-group $KK(A,B)$,
  • if $A=\mathbb{C}$, then we have $KK(\mathbb{C},B)\cong K_{0}(B)$, the operator $K$-theory group of $B$ in degree 00;
  • if $B=\mathbb{C}$, then we have $KK(A,\mathbb{C})\cong K^{0}(A)$, the $K$-homology of $A$ which consists of equivalence classes of $A$-Fredholm modules.
In order to define the $KK$-product, we need some special $C^{\ast}$-algebras.
@@
Kasparov showed that there is a product operation on $KK$-groups. We briefly explain how the $KK$-product is constructed.
Given $x\in KK(A,D)$, $y\in KK(D,B)$, choose representatives $(E_{1},\phi_{1},F_{1})\in\mathbb{E}(A,D)$ and $(E_{2},\phi_{2},F_{2})\in\mathbb{E}(D,B)$; then define $(E,\phi,F)\in\mathbb{E}(A,B)$ by $E=E_{1}\hat{\otimes}_{\phi_{2}}E_{2}$, $\phi=\phi_{1}\hat{\otimes}_{\phi_{2}}1$, $F=F_{1}\#F_{2}$, where $F_{1}\#F_{2}$ is a suitable combination of $F_{1}$ and $F_{2}$.
In general, $F$ will be a combination of $F_{1}\hat{\otimes}1$ and $1\hat{\otimes}F_{2}$ (when these are suitably defined). The coefficients will come from a judiciously chosen “partition of unity”. However, $1\hat{\otimes}F_{2}$ cannot be defined in general cases, so we need an operator $F\in\mathcal{L}(E)(E=E_{1}\hat{\otimes}_{\phi_{2}}E_{2})$ which “plays the role of $1\hat{\otimes}F_{2}$ up to compacts.”
For any $x\in E_{1}$ there is an operator $T_{x}\in\mathcal{L}(E_{2},E)$ defined by $T_{x}(y)=x\hat{\otimes}y$.
Definition 1.5.11.
  • An operator $F\in\mathcal{L}(E)$ is called an $F_{2}$-connection for $E_{1}$(or an $F_{2}$-connection on $E$) if, for any $x\in E_{1}$,
    $T_{x}\circ F_{2}-(-1)^{\partial x\cdot\partial F_{2}}F\circ T_{x}\in\mathcal{K}(E_{2},E),$
    $F_{2}\circ T_{x}^{\ast}-(-1)^{\partial x\cdot\partial F_{2}}T_{x}^{\ast}\circ F\in\mathcal{K}(E,E_{2}),$
  • Let $F$ be an $F_{2}$-connection on $E$. $(E,\phi,F)$ is called a Kasparov product for $(E_{1},\phi_{1},F_{1})$ and $(E_{2},\phi_{2},F_{2})$ if
    1. (1)
      $(E,\phi,F)$ is a Kasparov $(A,B)$-module;
    2. (2)
      For all $a\in A$, $\phi(a)[F_{1}\hat{\otimes}1,F]\phi(a)^{\ast}\geq 0\text{ mod }\mathcal{K}(E)$.
Let $G$ be any $F_{2}$-connection of degree $1$. Using Kasparov’s Technical Theorem (Theorem 1.4 in [kasparov1988equivariant]), we can choose appropriate $M,N\in\mathcal{L}(E)$ with $0\leq M,N\leq 1$ and $M+N=1$ satisfying several conditions so that the above (1) and (2) are realized, and the product operator $F$ can be constructed to be $F=M^{1/2}(F_{1}\hat{\otimes}1)+N^{1/2}G$,
@@
Let $A$, $B$ be separable and $D$ be $\sigma$-unital graded $C^{\ast}$-algebras. Let $(E_{1},\phi_{1},F_{1})$ be a Kasparov $(A,D)$-module and $(E_{2},\phi_{2},F_{2})$ be a Kasparov $(D,B)$-module, there is a Kasparov product for $(E_{1},\phi_{1},F_{1})$ and $(E_{2},\phi_{2},F_{2})$, which is a Kasparov $(A,B)$-module, unique up to homotopy $\sim_{h}$. The Kasparov product defines a bilinear function ${\otimes}_{D}:KK(A,D)\times KK(D,B)\rightarrow KK(A,B)$.
Remark 1.5.13.
Such a product between $KK$-groups does not exist in algebraic $K$-theory.
@@
In the discussions in Chapter 5, we also need the $\mathcal{R}KK$-group which is defined in [kasparov1988equivariant]. We give a brief introduction here.
Definition 1.5.15.
For a $\sigma$-compact $G$-space $M$ and two $G-C_{0}(M)$-algebras $A$ and $B$, define the group $\mathcal{R}KK^{G}(M;A,B)$ in the same way as $KK^{G}(A,B)$, with the additional condition on $(E,\phi,F)\in\mathbb{E}^{G}(A,B)$ that: for any $a\in A$, $b\in B$, $f\in C_{0}(M)$, $e\in E$, we have the equality $(fa)eb=ae(fb)$ in $E$.
@@ 1.6 Heisenberg calculus for pseudo-differential operators
The normalization of operators to 00-order and the inversion of cosymbol classes modulo smoothing terms are essential in our theory, but they are outside of the class of differential operators which we discussed in Section 1.4. In this section, we consider the Heisenberg calculus for pseudo-differential operators.
@@
A subset $X$ of a groupoid $\mathcal{G}$ is proper if both $r:X\rightarrow M$ and $s:X\rightarrow M$ are proper maps. Denote by $\Omega_{r}$ and $\Omega_{s}$ the bundle of smooth densities tangent to the range and source fibers of $\mathcal{G}$ respectively. The space of smooth sections of $\Omega_{r}$ whose support is a proper subset of $\mathcal{G}$ is denoted $C_{p}^{\infty}(\mathcal{G};\Omega_{r})$.
@@
In [connes1994noncommutative], Connes constructed the tangent groupoid
$\mathbb{T}M=TM\times\{0\}\sqcup(M\times M)\times\mathbb{R}^{\times},$
and used it to prove the Atiyah-Singer’s index theorem in $K$-theory. In the filtered case, a similar concept can be defined as follows.
@@
Define $\mathbb{G}r(H)$ to be the disjoint union of the osculating groupoid $Gr(H)$ with a family of pair groupoids indexed by ×=$\mathbb{R}^{\times}=\mathbb{R}\setminus\{0\}$,
@@
it is called the H-tangent groupoid of the osculating groupoid $Gr(H)$.
Theorem 1.6.5 ([van2017tangent]).
There is a compatible global smooth structure on $\mathbb{G}r(H)$ that makes $\mathbb{G}r(H)$ into a Lie groupoid.
@@
The cosymbol of a pseudo-differential operator can be viewed as a family of operators on the manifolds $Gr(H)_{x}$ for $x\in M$. So specifically, we need to analyze the special case $\mathcal{G}=Gr(H)$.
@@
Let $S^{k}(\mathcal{G})$ be the space of distributions $u\in\mathcal{E}^{\prime}(\mathcal{G})$ such that $u$ is transversal to the map $\pi:\mathcal{G}\to M$ and for each $\lambda\in\mathbb{R}^{+}$,
${\alpha_{\lambda}}_{\ast}u-\lambda^{k}u\in C_{c}^{\infty}(\mathcal{G}).$
Let $u\in S^{k}(\mathcal{G})$. For each $x\in M$, the operator can be defined as
@@
Let $u\in S^{k}(\mathcal{G})$, then
  • If $k\leq 0$, then $\overline{Op(u)}\in\mathcal{L}(C^{\ast}(\mathcal{G}))$;
  • If $k<0$, then $\overline{Op(u)}\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$;
  • If $k>0$, and $u$ satisfy the Rockland condition, then $\overline{Op(u)}$ is a regular operator.
If $P\in\mathcal{U}^{k}(\mathfrak{g})$ be a left-invariant differential operator on $\mathcal{G}$, then the distribution $(u,f):=Pf(e)$ belongs to $S^{k}(\mathcal{G})$. Since $P$ is left-invariant, $Op(u)=P$. Thus we can consider $\mathcal{U}^{k}(\mathfrak{g})\subseteq S^{k}(\mathcal{G})$.
For left-invariant differential operators on $\mathcal{G}$, the Rockland condition facilitates the consideration in $KK$-theory:
@@
$D=\overline{P}/(1+\overline{P}^{\ast}\overline{P})^{1/2}\in\mathcal{L}(C_{0}(M;E)\otimes_{C_{0}(M)}C^{\ast}(\mathcal{G}),C_{0}(M;F)\otimes_{C_{0}(M)}C^{\ast}(\mathcal{G}))$
@@
For any arbitrary H-elliptic operator $P$, the principal cosymbol $\sigma_{H}(P)$ represents a $KK$-element $[\sigma_{H}(P)]\in KK(C_{0}(M),C^{\ast}(Gr(H))$.
@@
$\Big{[}C_{0}(M;E_{0}\oplus E_{1})\otimes_{C_{0}(M)}C^{\ast}(\mathcal{G}),\begin{pmatrix}0&\overline{P}^{\ast}/(1+\overline{P}\overline{P}^{\ast})^{1/2}\\ \overline{P}/(1+\overline{P}^{\ast}\overline{P})^{1/2}&0\end{pmatrix}\Big{]}$
represents an element in $KK(C_{0}(M),C^{\ast}(\mathcal{G}))$. Apply this result to $\sigma_{H}(P)$, the family of left-invariant differential operators on $Gr(H)$, so it represents an element in $KK(C_{0}(M),$ $C^{\ast}(Gr(H))$. ∎
This framework can be extended to pseudo-differential operators. For a pseudo-differential operator $P$ on $M$, the principal cosymbol $\sigma_{H}(P)$ is an unbounded multiplier of the $C^{\ast}$-algebra $C^{\ast}(Gr(H))$, and the H-order of $P$ less or equal to 00 decide the boundedness or compactness of $P$ as follows:
@@
Let $P$ be a pseudo-differential operators of H-order $k$ on compact manifold $M$, then
@@
1.7 Index map for H-elliptic operators
Let $M$ be a compact manifold with the filtration structure $H^{\bullet}$ and $E,F$ vector bundles over $M$, a sequence of Sobolev spaces $\tilde{H}^{s}(M,E,F)$ are defined in [androulidakis2022pseudodifferential]. For a non-negative integer $s$, it encompasses the set of vector bundle $E$-valued distributions $u$ such that $Du\in L^{2}(F)$ for any differential operators $D\in\text{Diff}^{s}_{{H}}(M,E,F)$. It can be naturally extended to real numbers $s\geq 0$, and for $s<0$, $\tilde{H}^{s}(M,E,F)$ is defined to be the dual of $\tilde{H}^{-s}(M,E,F)$.